# Reinforced Generation of Combinatorial Structures: Applications to Complexity Theory

### v4

@@ Abstract
We explore whether techniques from AI can help discover new combinatorial structures that improve on known limits on efficient algorithms. Specifically, we use AlphaEvolve (an LLM coding agent) to study two settings:
a) Average-case hardness for MAX-CUT and MAX-Independent Set: We improve a recent result of Kunisky and Yu to obtain near-optimal upper and (conditional) lower bounds on certification algorithms for MAX-CUT and MAX-Independent Set on random 3- and 4-regular graphs. Our improved lower bounds are obtained by constructing nearly extremal Ramanujan graphs on as many as $163$ nodes, using AlphaEvolve. Additionally, via analytical arguments we strengthen the upper bounds to settle the computational hardness of these questions up to an error in the third decimal place.
b) Worst-case Hardness of Approximation for MAX-k-CUT: We obtain new inapproximability results, proving that it is NP-hard to approximate $\textup{MAX-}{4}\textup{-CUT}$ and $\textup{MAX-}{3}\textup{-CUT}$ within factors of $0.987$ and $0.9649$ respectively, using AlphaEvolve to discover new gadget reductions. Our $\textup{MAX-}{4}\textup{-CUT}$ result improves upon the SOTA of $0.9883$, and our $\textup{MAX-}{3}\textup{-CUT}$ result improves on the current best gadget-based inapproximability result of $0.9853$, but falls short of improving the SOTA of $16/17$ that relies on a custom PCP, rather than a gadget reduction from “standard” Håstad-style PCPs.
A key technical challenge we faced: verifying a candidate construction produced by AlphaEvolve is costly (often requiring exponential time). In both settings above, our results were enabled by using AlphaEvolve itself to evolve the verification procedure to be faster (sometimes by $10,000\times$). We conclude with a discussion of norms by which to assess the assistance from AI in developing proofs.
@@ 1 Introduction
In this paper, we address the following question: Can AI based methods help us make novel and non-trivial discoveries in complexity theory? We provide evidence towards answering this in the affirmative, via progress on two problems.
@@
    Hardness of certifying upper bounds on the MAX-CUT and MAX-IND-SET of sparse random graphs: We improve upon the combinatorial structures presented in [26], achieving improved lower bounds in the $\{3,4\}$-regular cases. Together with new non-AI based algorithmic improvements [^1] These algorithmic improvements are derived from a discussion with Sidhanth Mohanty, Northwestern University., we achieve an almost tight resolution of these questions [^2] By almost tight, we mean that the lower bounds and upper bounds on efficient certificates match up to the second decimal place..
@@
    NP-Hardness of approximation for $\textup{MAX-}{k}\textup{-CUT}$. We find new gadget-based reductions (from standard Håstad PCPs [19]) for the NP-hardness of approximating $\textup{MAX-}{k}\textup{-CUT}$ for $k\in\{3,4\}$. For $\textup{MAX-}{4}\textup{-CUT}$, we improve on the current best inapproximability result from $0.9883$ [2] to $0.987$. For $\textup{MAX-}{3}\textup{-CUT}$, we improve on the current best gadget-based inapproximability result from $0.9853$ [24] to $0.9649$. Our result for $\textup{MAX-}{3}\textup{-CUT}$ is sandwiched between two results using custom PCPs for this problem: [14] with $0.9696$, and [2] with $0.9411$.
To this end we invoke AlphaEvolve [30], a coding agent for scientific and algorithmic discovery, to find and verify finite combinatorial structures that improve on prior constructions. At a high level AlphaEvolve is a tool that uses an LLM (Large-Language-Model) to iteratively evolve code snippets that generate combinatorial structures (we will sometimes refer to these as constructions) of high quality by some criterion. It is worth emphasizing that while the structures we generate using AlphaEvolve are finite constructions, our main results (Theorems 3.1 and 4.1) embed the universal quantifier $\forall n$, with $n$ being the size of the problem instance, thanks to appropriate “lifting” [26, 36] arguments.
Operationally, the system consists of three parts: a) a code snippet ($\mathbf{C}$) that constructs combinatorial structures, b) An evaluation function (referred to as the verifier) that verifies and scores the generated structures, and c) An LLM that suggests a new search function $\mathbf{C}_{\texttt{new}}$ based on the set of prior $\mathbf{C}$’s and the history of constructions. Via prompting the LLM, the objective is to nudge $\mathbf{C}_{\texttt{new}}$ towards better structures. (Section 2 provides a more detailed overview of AlphaEvolve.)
@@
Speeding up verification in AlphaEvolve: The constructions we search for in both problems are undirected graphs. For the hardness of certifying upper bounds for sparse random graphs, we use AlphaEvolve to search for $d=\{3,4\}$-regular Ramanujan graphs, along with a witness for a large cut/independent set. Evaluating whether a graph is Ramanujan, and whether or not the cut/independent set is a valid witness, are efficient. As a result, for this problem we could find and verify graphs with as many as $163$ nodes.
@@
However in studying the NP-Hardness of approximation for $\textup{MAX-}{k}\textup{-CUT}$, we seek gadgets that reduce the 3LIN(k) problem to MAX-k-CUT; here the verification is unfortunately not efficient: it requires checking the gadget against $\Omega(k^{m})$ linear constraints, where $m$ is the total number of variables in the target problem. A standard methodology in gadget-based hardness reductions is to add auxiliary variables to have flexibility in mapping the source and the target problems [36]. This exacerbates the exponential running time of the verifier (for reference, our final gadgets contain $m=14$ and $m=19$ variables for $k=3$ and $k=4$ respectively). To address this we invoke AlphaEvolve to speed up the verifier, as follows. (In Section 4.1 we discuss this in more detail.)
We began with a code snippet containing a brute-force implementation of the verifier, using AlphaEvolve to iteratively improve its runtime. The main challenge we faced was ensuring that the evolved snippets were correct, in the sense of being functionally equivalent to the original verifier. We found that the following two ideas sufficed to achieve large speedups while maintaining correctness.
  • We used prior code snippets $\mathbf{C}_{\texttt{old}}$, whose correctness has already been verified, to evaluate synthetically generated gadgets. We compare the outputs of a new candidate $\mathbf{C}_{\texttt{new}}$ to this dataset to ensure that it is also correct.
  • We used a separate LLM as a judge to ensure that all candidates $\mathbf{C}_{\texttt{new}}$ are correct implementations of the verifier.
@@
We prompted AlphaEvolve to improve the execution time of the verifier, while ensuring that it passes the two correctness tests above. To our surprise, AlphaEvolve improved the running time by $10,000\times$ for our final $\textup{MAX-}{4}\textup{-CUT}$ gadget of size $m=19$, and by a similar factor for our final $\textup{MAX-}{3}\textup{-CUT}$ gadget of size $m=14$. Furthermore, we found the improved versions to be provably correct by human inspection (given there was a possibility of being correct only on the synthetic dataset). The improvement came from a number of system-level improvements together with a branch-and-bound strategy to prune many of the $\Omega(k^{m})$ constraints.

### v5

# Reinforced Generation of Combinatorial Structures: Hardness of Approximation

@@ Abstract
Can AI based methods help us make advances in complexity theory? We provide evidence towards answering this in the affirmative, using AlphaEvolve (an LLM code mutation agent) to obtain new results in three settings:
a) Average-case hardness for MAX-CUT and MAX-Independent Set: We improve a recent result of Kunisky and Yu to obtain near-optimal upper and (conditional) lower bounds on certification algorithms for MAX-CUT and MAX-Independent Set on random 3- and 4-regular graphs. Our improved lower bounds are obtained by constructing nearly extremal Ramanujan graphs on as many as $163$ vertices. Additionally, via analytical arguments we strengthen the upper bounds to settle the computational hardness of these questions up to an error in the third decimal place.
b) Worst-case Hardness of Approximation for MAX-k-CUT: We obtain new inapproximability results, proving that it is NP-hard to approximate $\textup{MAX-}{4}\textup{-CUT}$ and $\textup{MAX-}{3}\textup{-CUT}$ within factors of $0.987$ and $0.9649$ respectively, using AlphaEvolve to discover new gadget reductions. Our $\textup{MAX-}{4}\textup{-CUT}$ result improves upon the SOTA of $0.9883$, and our $\textup{MAX-}{3}\textup{-CUT}$ result improves on the current best gadget-based inapproximability result of $0.9853$, but falls short of improving the SOTA of $16/17$ that relies on a custom PCP (rather than a gadget reduction from “standard” Håstad-style PCPs).
c) Worst-case Hardness of Approximation for the metric Traveling Salesman Problem (TSP): We show that it is NP-hard to approximate the minimum cost tour within a factor of $111/110$ using AlphaEvolve to discover a new gadget, thus improving the SOTA of $117/116$. We provide a modular soundness and completeness argument for the reduction of $3\textup{LIN}({2})$ (a standard constraint satisfaction problem used in hardness reductions) to TSP, which enabled AlphaEvolve to search over finite constraint graphs. This modularization may be of independent interest for future work on TSP inapproximability.
A key technical challenge we faced: verifying a candidate construction produced by AlphaEvolve is costly (sometimes requiring time exponential in the size of the construction). Our results were enabled by using AlphaEvolve itself to evolve the verification procedure to be faster (sometimes by $10,000\times$ for our gadgets). Our results suggest that gadget based proofs would benefit from a pass through AI-based tools to obtain stronger results.
@@ 1 Introduction
In this paper, we study the following question: Can AI based methods help us make novel and non-trivial discoveries in complexity theory? We provide evidence towards answering this in the affirmative, via progress on three problems.
@@
    Hardness of certifying upper bounds on the MAX-CUT and MAX-IND-SET of sparse random graphs. In Section 2, we improve upon the combinatorial structures presented in [36], achieving improved lower bounds in the $\{3,4\}$-regular cases. Complementing this result, we obtain new upper bounds which improve upon Hoffman’s classic spectral bounds [30, 25]. (These upper bounds are derived via analytical approaches [^1] These improvements are derived from a discussion with Sidhanth Mohanty, Northwestern University..) Together, these achieve an almost tight resolution to the problem. [^2] By this we mean: the lower bounds and upper bounds on efficient certification algorithms match up to the second decimal place..
@@
    NP-Hardness of approximation for $\textup{MAX-}{k}\textup{-CUT}$. In Section 3, we give new gadget-based reductions (from standard Håstad PCPs [26]) for the NP-hardness of approximating $\textup{MAX-}{k}\textup{-CUT}$ for $k\in\{3,4\}$. For $\textup{MAX-}{4}\textup{-CUT}$, we improve on the current best inapproximability result from $0.9883$ [3] to $0.987$. For $\textup{MAX-}{3}\textup{-CUT}$, we improve on the current best gadget-based inapproximability result from $0.9853$ [32] to $0.9649$. Our result for $\textup{MAX-}{3}\textup{-CUT}$ is sandwiched between two results using custom PCPs for this problem: [21] with $0.9696$, and [3] with $0.9411$.
  • c)
    NP-Hardness of approximation for metric TSP. In Section 4, we give a new gadget based reduction (from standard Håstad PCPs [26]) for the NP-hardness of approximating metric TSP. We improve the current SOTA of $117/116$ [11] to $111/110$. Our improvement advances the long line of work establishing hardness results for this problem [45, $1+\delta$] (without an explicit value for $\delta$), [16, 5381/53805381/5380][7, 3183/31823183/3182][44, 220/219220/219][37, 185/184185/184][33, 123/122123/122][11, $117/116$]. As an aside, our AI-based exploration is ongoing, with the potential for further improvements to this bound.
Notably, our results come from a single (AI-derived) technique, AlphaEvolve [42, 48], to find and verify finite combinatorial structures that improve on prior constructions. At a high level AlphaEvolve uses an LLM (Large-Language-Model) to iteratively evolve code snippets that generate combinatorial structures (we will sometimes refer to these as constructions) that are of high quality by some criterion. While the structures we generate using AlphaEvolve are finite constructions, our main results (Theorems 2.13.1, and 4.1) embed the universal quantifier $\forall n$, with $n$ being the size of the problem instance, thanks to appropriate “lifting” [36, 50, 11] arguments. At a minimum our work suggests that we would do well to routinely run gadget proofs through an “AI optimization” phase. We present our results in increasing order of human involvement (required to make the problem amenable to AlphaEvolve).
Operationally, the system consists of three parts: a) a code snippet ($\mathcal{C}$) that constructs combinatorial structures, b) An evaluation function (referred to as the verifier) that verifies and scores the generated structures, and c) An LLM that suggests a new code snippet $\mathcal{C}_{\texttt{new}}$ based on the set of prior $\mathcal{C}$’s and the history of constructions. Via prompting the LLM, the objective is to nudge $\mathcal{C}_{\texttt{new}}$ towards building better structures. (Appendix A provides a more detailed overview of AlphaEvolve.)
@@
Speeding up verification in AlphaEvolve: The combinatorial structures sought in these problems are based on undirected graphs. While brute-force verification of our structures was reasonably fast for the TSP lower bound and for certifying upper bounds on sparse random graphs, it was significantly more challenging for $\textup{MAX-}{k}\textup{-CUT}$. We discuss this in more detail below.
For TSP, one of the discovered gadgets had $20$ edges (over $12$ vertices) along with a large number of constraints ($\approx 11!$) for a successful verification. As a result, we needed some care in writing a speedy verifier that operated within roughly one second.
@@
For the hardness of certifying upper bounds for sparse random graphs, we use AlphaEvolve to search for $d=\{3,4\}$-regular Ramanujan graphs, along with a witness for a large cut/independent set. Evaluating whether a graph is Ramanujan, and whether or not the cut/independent set is a valid witness, are efficient. As a result, for these problems we could find and verify graphs with as many as $163$ vertices.
@@
In studying the NP-Hardness of approximation for $\textup{MAX-}{k}\textup{-CUT}$, we seek gadgets that reduce the $3\textup{LIN}({k})$ problem [26] to MAX-k-CUT ($3\textup{LIN}({k})$ corresponds to a set of constraints of the form $x+y+z\equiv b\pmod{k}$, where $x,y,z\in[0,\ldots,k-1]$). Here the verification is unfortunately not efficient. It requires checking the gadget against $\Omega(k^{m})$ linear constraints, where $m$ is the total number of variables in the target problem. A standard technique in gadget-based hardness reductions is to add auxiliary variables to have flexibility in mapping the source and the target problems [50]. This exacerbates the exponential running time of the verifier (for reference, our final gadgets contain $m=14$ and $m=19$ variables for $k=3$ and $k=4$ respectively). To address this we invoke AlphaEvolve to speed up the verifier, as follows.
@@
We prompted AlphaEvolve to improve the execution time of the verifier, while ensuring that it passes various correctness checks. To our surprise, AlphaEvolve improved the running time by $10,000\times$ for our final $\textup{MAX-}{4}\textup{-CUT}$ gadget of size $m=19$, and by a similar factor for our final $\textup{MAX-}{3}\textup{-CUT}$ gadget of size $m=14$. We used human inspection to verify the correctness of these improved verifiers (given their possibility of being correct only on synthetically generated gadgets used for correctness checks). The improvement came from a number of system-level improvements together with a branch-and-bound strategy to prune many of the $\Omega(k^{m})$ constraints. We discuss this in more detail in Section 3.2.